if abc are in ap then e^9ax+k),e^(bx+k),e^(cx+k) are in

Grade:12th pass

1 Answers

Arun
6 years ago
 
If we take ratio like,
e^(bx+ k)/ e^(ax +k)   = e^(cx +k)/ e^(bx+ k)
Now we get
e^(b-a)x  =  e^(c-b)x
When base is same, then power is compared 
Hence
b-a = c- b
2b = a+c
This is the given condition.
Hence the terms will be in G.P.